cost of 27 toasters and 18 mixers is rs. 80100.what is the cost of 12 toasters and 8 mixers together?
Answer by Alan3354(69443)

cost of 27 toasters and 18 mixers is rs. 80100.what is the cost of 12 toasters and 8 mixers together?
---------
12/27 = 4/9
8/18 = 4/9
--------------
--> 4/9 as much